I need help solving this Quadratic equation: x+11√ ̅x +10.
I know how to solve a simple trinomial but the radical sgn confuses me.

Answer by nerdybill(7384)

I'm assuming you meant the equation to equal to zero:
x+11√ ̅x +10 = 0
.
Think of this as a "quadratic".
Let y = √ ̅x
then
y^2 = x
.
so, we can rewrite original equation as:
y^2 + 11y + 10 = 0
(y+10)(y+1)=0
y = {-10, -1}
.
since
x = y^2
our solution is then
x = {100, 1}
we can throw out the 100 (extraneous) leaving:
x = 1